OpenStudy (anonymous):

y is less than or equal to 2(x-3)^2-2

OpenStudy (anonymous):

\[y \le 2(x-3)^{2}-2\]

OpenStudy (anonymous):

Have to solve for x's

OpenStudy (anonymous):

\[ 2 (x-3)^2-2=2 x^2-12 x+16=2 (x-4) (x-2) \]

OpenStudy (anonymous):

Can you finish it now?

OpenStudy (agent0smith):

You can graph this, it's an inequality with two variables. There are infinite solutions (hence why you graph it).

OpenStudy (anonymous):

Thanks! I think so...